tyler is 20 years older than shane. in 5 years, shane will be half as old as tyler is then. use 2 variables to find their present ages.
Answer by stanbon(75887)

tyler is 20 years older than shane. in 5 years, shane will be half as old as tyler is then. use 2 variables to find their present ages.
-------
Equations:
t = s + 20
s+5 = (1/2)(t+5)
-----------------------
Substitute for "t" and solve for "s":
s+5 = (1/2)(s+25)
---
2s + 10 = s+25
s = 15 (Shane's age now)
-----
Solve for "t":
t = s + 20
t = 35 (Tyler's age now)
===============
Cheers,
Stan H.
==================